OpenStudy (anonymous):

What is the solution to the equation 9(x + 1) = 27?

OpenStudy (anonymous):

og(9^(x - 1)) = log(27) (x - 1)*log(9) = log(27) x - 1 = log(27) / log(9) x - 1 = 1.5 x = 1.5 + 1 x = 2.5?

OpenStudy (swissgirl):

Ok first step is to distribute the 9 9x+9=27 Then subtract or add any constants. IN this case subtract the 9 from both sides 9x+9-9=27-9 9x=18 Now divide 9 from both sides 9x/9=18/9 x=2

OpenStudy (anonymous):

Please don't break the Code of Conduct by giving out answers. @swissgirl

OpenStudy (swissgirl):

OHHHH NOOOO :O IRESPECT the rules and Code of Conduct given to the OS Society

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

Excuse me? How old are you?

OpenStudy (anonymous):

Childish - I can see.

OpenStudy (swissgirl):

lol im only 10 dont expect much from me

OpenStudy (anonymous):

I wouldn't be surprised if you actually were.

OpenStudy (swissgirl):

Miss Respect pls dont take this as a personal offence. I meant no harm and I choose to abide by the rule sand regulations provided by the OS legislators

OpenStudy (anonymous):

Then please abide by the Code of Conduct by not giving out answers, provide an explanation instead.

OpenStudy (swissgirl):

sure dear ... keep up the good work The world needs more good doers such as yourself

OpenStudy (anonymous):

lol both up y'all give it up, move on

OpenStudy (swissgirl):

hahaha

OpenStudy (anonymous):

LOL I love it when people try be snitches....